OpenStudy (anonymous):

Use basic identities to simplify the expression. cos θ - cos θ sin^2θ

OpenStudy (anonymous):

First you could factor out cos(theta)

OpenStudy (anonymous):

\[\Large \cos(\theta)(1-\sin^2(\theta))\]

OpenStudy (anonymous):

Can you go from there?

OpenStudy (anonymous):

How did you factor out cos (theta)? Did you divide by cos(theta)?

OpenStudy (isaiah.feynman):

@danber8 your final answer would be \[\cos ^{3}\]

OpenStudy (isaiah.feynman):

cos^3 theta

OpenStudy (anonymous):

Iseah, I'm confused as to how you got that answer

OpenStudy (anonymous):

You don't divide, you see that therer's cos(theta) being multipled by those two terms so you FACTOR it not divide, factoring doesn't change the value

OpenStudy (anonymous):

@Isaiah.Feynman thats why you're discouraged to give answers it doesn't help in the learning process, anyways \[\Large \cos(\theta)(1-\sin^2(\theta))=\cos(\theta)-\cos(\theta)\sin^2(\theta)\] your original equation, all we did was factor out a cos(theta)

OpenStudy (anonymous):

I understand that now

OpenStudy (anonymous):

Ok and do you have you're list of trig identities?

OpenStudy (anonymous):

u mean like cos (theta) =sin(pi/2 - theta)?

OpenStudy (anonymous):

Yes precisly and one of them is \[\sin^2(\theta)+\cos^2(\theta)=1\] which can be moved around to \[\cos^2(\theta)=1-\sin^2()\]

OpenStudy (anonymous):

\[\cos^2(\theta)=1-\sin^2(\theta)*\]

OpenStudy (anonymous):

so now you can replace cos^2(theta) in the parentheses

OpenStudy (anonymous):

Would we divide both sides by cos (theta)?

OpenStudy (anonymous):

\[\cos(\theta)(1-\sin^2(\theta))=\cos(\theta)(\cos^2(\theta))\]

OpenStudy (anonymous):

We're not dividing by anything, we don't even this equal to something, its jsut and expression like 3x+5, its not being set to anything all they're asking is to simplify

OpenStudy (anonymous):

How did you know cos(theta)(1-sin^2(theta))= cos(theta)(cos^2(theta))?

OpenStudy (anonymous):

because one of our trig identities is \[\Large \sin^2(\theta)+\cos^2(\theta)=1\] which when we subtract \[\Huge \sin ^{2}\left( \theta \right)\] form both sides we get \[\Large \cos^2(\theta)=1-\sin^2(\theta)\]

OpenStudy (anonymous):

so now plug the left side in our problem

OpenStudy (anonymous):

So cos(theta) x cos^2(theta)=cos^3(theta)?

OpenStudy (anonymous):

costheta-costhetasin^2theta costheta(1-sin^2theta) costheta(cos^2theta) cos^3theta

OpenStudy (anonymous):

So far after factoring we were at cos(x)(1-sin^2(x)) now since we idntifies 1-sin^2(x) to be cos^2(x) so \[\Large \cos(\theta)(1-\sin^2(\theta))=\cos(\theta)(\cos^2(\theta))=\cos^3(\theta)\]

OpenStudy (anonymous):

I understand now, thanks

OpenStudy (anonymous):

Np.